# spacing=nonfrench

### v1

@@ 4.2.1 On a Theorem of Murai
If $\Gamma$ is a Lipschitz curve with Lipschitz constant $M$ then we have, for $0<s<1$,
@@ 5.1 The Conjugate Operator
Let $\Gamma$ be a chord-arc curve of length $2\pi$, and $1<p<\infty$, $0<s<1$. We have seen two ways of generalizing the fractional Sobolev spaces from $\mathbb{R}$ (or from $\mathbb{T}$) to $\Gamma$. The first one, which can be called the Douglas way is
$B_{p,p}^{s}(\Gamma)=\bigg\{f\in L^{p}(\Gamma):\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty\bigg\}.$
The second one, via Littlewood-Paley theory, is
$\mathcal{B}_{p,p}^{s}(\Gamma)=\bigg\{f\in L^{p}(\Gamma):\iint_{\Omega_{i,e}}|\nabla u_{i,e}(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<\infty\bigg\}.$
Here, $u_{i,e}$ stands for the harmonic extension of $f$ to $\Omega_{i,e}$. We have seen that Plemelj-Calderón problem is solvable in some sense on both spaces $B_{p,p}^{s}(\Gamma)$ and $\mathcal{B}_{p,p}^{s}(\Gamma)$ so that the natural question arises of whether these spaces coincide.
Assume that $\Omega$ is bounded and $\mathbb{D}$ is the unit disk. Let $z_{0}$ be a point in $\Omega$. If $u$ is harmonic in $\Omega$ it is well known that there exists a unique harmonic function $\tilde{u}$ in $\Omega$ such that $\tilde{u}(z_{0})=0$ and $u+i\tilde{u}$ is holomorphic in $\Omega$. In the case of $\Omega=\mathbb{D}$, $\tilde{u}$ is the harmonic (Poisson) extension of $H(f)$ where $H$ is the Hilbert transform:
@@
By Cauchy-Riemann equations, $|\nabla\tilde{u}|=|\nabla u|$, so that $\mathcal{B}_{p,p}^{s}(\Omega)$ is conjugate-invariant. We denote the subspace of $\mathcal{B}_{p,p}^{s}(\Omega)$ consisting of holomorphic functions by $\mathcal{HB}_{p,p}^{s}(\Omega)$, i.e., the set of $u+i\tilde{u}$ with $u,\tilde{u}\in\mathcal{B}_{p,p}^{s}(\Omega)$.
A necessary condition for $B_{p,p}^{s}(\Gamma)=\mathcal{B}_{p,p}^{s}(\Gamma)$ to hold is thus that $B_{p,p}^{s}(\Gamma)$ is stable by conjugation. Let $\varphi:\,\mathbb{D}\to\Omega$ be the Riemann map with $\varphi(0)=z_{0}$. Recall that $t\mapsto z(t)$ denotes the arc-length parametrization of $\Gamma$ and $\lambda(e^{it})=z(t)$ so that $\varphi=\lambda\circ h$ where $h$ is an absolutely continuous homeomorphism of $\mathbb{T}$ with $|h^{\prime}|=|\varphi^{\prime}|$. Recall that $\lambda$ is a bi-Lipschitz homeomorphism from $\mathbb{T}$ onto $\Gamma$. A simple computation shows that if $f\in B_{p,p}^{s}(\Gamma)$; that is $g=f\circ\lambda\in B_{p,p}^{s}(\mathbb{T})$, and $u$ is its harmonic extension to $\Omega$, then $\tilde{u}$ is the harmonic extension of $\tilde{f}=\tilde{g}\circ\lambda^{-1}$. Here,
@@
It is known (see [9], p.247 in [16]) that $V_{h}^{-1}HV_{h}$ is bounded on $L^{p}(\mathbb{T})$ if and only if $|h^{\prime}|=|\varphi^{\prime}|$ belongs to $A_{p}$ on $\mathbb{T}$. Based on this comment on the limiting case $s=0$, we now state a theorem about the general case $0\leq s\leq 1$.
@@
For $p>1,\,0\leq s\leq 1$, suppose $\Gamma$ is such that
@@
Then, $B_{p,p}^{s}(\Gamma)$ is stable by conjugation, and moreover, the conjugate operator $f\mapsto\tilde{f}$ on $B_{p,p}^{s}(\Gamma)$ is bounded.
Before going to the proof, let us comment on the hypothesis of Theorem 5.1. If $\omega$ is an $A_{\infty}$ weight let us define $\delta(\omega)=\inf\{q>1:\,\omega\in A_{q}\}$. It is known [9] that if $\omega\in A_{q}$ for some $q>1$ then $\delta(\omega)<q$ (see p.254 in [16]). Let us then define $\delta=\delta(|\varphi^{\prime}|)$. Noting that the hypothesis imply that $|\varphi^{\prime}|\in A_{2}$, we see that the hypothesis of the theorem are equivalent to
$\delta<q<\frac{\delta}{\delta-1}.$
On the other hand, it is known (see [21]) that $\Gamma$ being chord-arc implies $|\varphi^{\prime}|$ having $A_{\infty}$ but there are examples of chord-arc curves such that $|\varphi^{\prime}|\notin A_{2}$ (see [23]), we thus know the chord-arc condition is not sufficient for Theorem 5.1 to hold.
@@
Specially, a natural condition on the curve $\Gamma$ implying that $|\varphi^{\prime}|\in A_{2}$ is the radial-Lipschitz condition. Indeed, if $\Gamma$ is a radial-Lipschitz curve it can be shown that $|\varphi^{\prime}|$ satisfies the Helson-Szegö condition (see section 5.3 for a specific proof): $\log|\varphi^{\prime}|=u+Hv$ with $u\in L^{\infty},\;v\in L^{\infty}$ and $\|v\|_{\infty}<\pi/2$, which is equivalent to $|\varphi^{\prime}|\in A_{2}$ as follows from the Helson-Szegö theorem ([16]). We may now state:
@@
For $0\leq s\leq 1$, if $\Gamma$ is such that $|\varphi^{\prime}|\in A_{2}$ then $B_{2,2}^{s}(\Gamma)$ is stable by conjugation, and moreover, the conjugate operator $f\mapsto\tilde{f}$ on $B_{2,2}^{s}(\Gamma)$ is bounded. The conclusion holds in particular for radial-Lipschitz curves.
For $0\leq s\leq 1$, $1<p<\infty$, we denote the subspace of $B_{p,p}^{s}(\Gamma)$ consisting of functions $f+i\tilde{f}$ such that $f,\;\tilde{f}\in B_{p,p}^{s}(\Gamma)$ by $\text{HB}_{p,p}^{s}(\Gamma)$.
@@ 5.2 The Operator $V_{s,p}$ on $\mathcal{B}^{s}_{p,p}(\Omega)$
Recall that $\mathcal{B}^{s}_{p,p}(\Omega)$, $0\leq s<1$ and $1<p<\infty$, is the set of harmonic functions $u$ on the domain $\Omega$ bounded by the Jordan curve $\Gamma$ such that its norm $\|u\|_{\mathcal{B}^{s}_{p,p}(\Omega)}<\infty$. Suppose $\Gamma$ is chord-arc which in particular is rectifiable. It is known that the non-tangential boundary values $f$ of $u$ exist almost everywhere on $\Gamma$ with respect to the arc-length measure and $u$ can be recovered from $f$ using the harmonic extension. The space of boundary functions of $\mathcal{B}^{s}_{p,p}(\Omega)$ is denoted by $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)$ so that $\mathcal{B}_{p,p}^{s}(\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)\cap\mathcal{B}_{p,p}^{s}(\Omega_{e}\!\to\!\Gamma)$. We may identify $\mathcal{B}_{p,p}^{s}(\Omega)$ with $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)$ for convenience, and will switch between the two freely. Similarly, for $0\leq s\leq 1$ and $1<p<\infty$ we also identify $B_{p,p}^{s}(\Gamma)$ with the set $B_{p,p}^{s}(\Omega)$ of its harmonic extensions in $\Omega$.
Let $\Omega,\,\Omega^{\prime}$ be two Jordan domains containing 0 and $\varphi$ a holomorphic diffeomorphism from $\Omega$ onto $\Omega^{\prime}$ fixing 0. Let $f\in\mathcal{HB}_{p,p}^{s}(\Omega^{\prime})$. Using the change of variable $\zeta=\varphi(z)$ we have
@@
By the Koebe distortion theorem (2.1) there exists a universal constant $C>1$ such that
@@
Notice that for s=11/ps=1-1/p, this operator is nothing but the composition by $\varphi$ and that, in this case, $V_{1-1/p,p}$ is an isometry.
For convenience of later use, when $p=2$, the operator $V_{s,p}$ will be simplified to be $V_{s}$.
5.2.1 The Operator $V_{s}$ on $\mathcal{B}^{s}_{2,2}(\Omega)$
In order to understand better the operator $V_{s}$ let us rewrite it by using an integration by parts:
@@ 5.2.1 The Operator $V_{s}$ on $\mathcal{B}^{s}_{2,2}(\Omega)$
where the supremum is taken over all sub-arcs $I$ of $\mathbb{T}$ and $g_{I}$ denotes the integral mean of $g$ over $I$. It is known that $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$ if $\varphi(\mathbb{D})$ is chord-arc, but not vice versa. By Fefferman-Stein, $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$ if and only if
@@
is a Carleson measure in $\mathbb{D}$ (see e.g. Ch.VI in [16]).
@@
Suppose $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$; that is, $\mu$ is a Carleson measure in $\mathbb{D}$. Let $f\in E^{2}(\Omega)$ so that $T_{0}f\in E^{2}(\mathbb{D})$. Then by Carleson (see e.g. Theorem 3.9 in [16])
@@
Combined with (5.2), this completes the proof of statement $(1)$.
@@
or, in other words, that $\mathsf{g}(f\circ\varphi)\in L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$. Since $\Gamma$ is assumed to be chord-arc, we have that $|\varphi^{\prime}|$ has the weight $A_{\infty}$ on $\mathbb{T}$. We can then apply a theorem of Gundy-Wheeden [19] (see also [21]) which implies that $\mathsf{g}(f\circ\varphi)\in L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$ if and only if the non-tangential maximal function $\mathsf{n}(f\circ\varphi)$ of $f\circ\varphi$ is in $L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$. On the other hand, for any "circular curves" $\Gamma_{r}$ it holds that (see p.233 in [21])
@@
Notice that for this theorem we do not need $|\varphi^{\prime}|$ having $A_{2}$ on $\mathbb{T}$. This condition is nevertheless necessary for the following corollary to hold by Corollary 5.2 since if it is not attached then the (real) space of real parts of $E^{2}(\Omega\!\to\!\Gamma)$-functions is a proper subspace of the real space $L_{\mathbb{R}}^{2}(\Gamma,d\sigma)$
@@
Let $\Omega$ be a chord-arc domain bounded by $\Gamma$. If $|\varphi^{\prime}|\in A_{2}$ then $B_{2,2}^{0}(\Gamma)=\mathcal{B}_{2,2}^{0}(\Omega\!\to\!\Gamma)$.
5.3 The Equality $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$
The aim of this part is to prove the following theorem that is one of the main theorems of this paper.

### v2

@@ 4.3 Plemelj-Calderón problem on Lipschitz curves
If $\Gamma$ is a Lipschitz curve with Lipschitz constant $M$, then we have, for $0<s<1$,
@@ 5 Douglas versus Littlewood-Paley: the chord-arc case
Let $\Gamma$ be a chord-arc curve of length $2\pi$ with 0 in its inner domain $\Omega$, and $1<p<\infty$, $0<s<1$. Recall that we have two ways of generalizing fractional Besov-Sobolev spaces from the unit circle $\mathbb{T}$ to $\Gamma$. The first one can be called the Douglas way: we define $B_{p,p}^{s}(\Gamma)$ as the set of functions $f\in L^{p}(\Gamma)$ such that
$\|f\|_{B_{p,p}^{s}(\Gamma)}^{p}:=\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty,$
which are intermediate spaces between $B_{p,p}^{0}(\Gamma)$ (i.e., $L^{p}(\Gamma)$) and $B_{p,p}^{1}(\Gamma)$.
Let $B_{p,p}^{s}(\Omega)$, $0\leq s\leq 1$, be the set of harmonic extensions $u$ in $\Omega$ of all $f\in B_{p,p}^{s}(\Gamma)$. Then $B_{p,p}^{0}(\Omega)=E^{p}(\Omega)\oplus\overline{E^{p}(\Omega)}$. More precisely, there exists a unique pair of holomorphic functions $\varphi$ and $\psi$ in $E^{p}(\Omega)$ with $\varphi(0)=u(0)$ and $\psi(0)=0$ such that $u=\varphi+\overline{\psi}$. Since $\Gamma$ is rectifiable, it makes sense to speak of a tangential direction almost everywhere, and each function $\varphi$ in $E^{p}(\Omega)$ has a non-tangential limit almost everywhere on $\Gamma$, and conversely, $\varphi$ is the harmonic extension of its boundary limit function. Furthermore, since $\Omega$ is a chord-arc domain, in particular a Smirnov domain, $E^{p}(\Omega)$ coincides with the $L^{p}(\Gamma)$ closure of polynomials. Based on this, $E^{p}(\Omega)$ can be identified with the set of its boundary functions, so it is reasonable to identify $B_{p,p}^{0}(\Omega)$ and $B_{p,p}^{0}(\Gamma)$. Finally, for any $0\leq s\leq 1$, since the spaces $B_{p,p}^{s}(\Omega)$ decrease with $s$, we can identify $B_{p,p}^{s}(\Omega)$ and $B_{p,p}^{s}(\Gamma)$ for convenience and switch between the two freely.
The second one is defined by Littlewood-Paley theory: consider the set of functions $f\in L^{p}(\Gamma)$ such that its harmonic extension $u$ in $\Omega$ satisfies
$\|u\|_{\mathcal{B}_{p,p}^{s}(\Omega)}^{p}:=\iint_{\Omega}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<\infty.$
The set of such functions $u$ is denoted by $\mathcal{B}_{p,p}^{s}(\Omega)$.
In this section, our main goal is to prove $B_{p,p}^{s}(\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega)$ under some condition on $\Gamma$, or, more precisely, $\|f\|_{B_{p,p}^{s}(\Gamma)}\simeq\|u\|_{\mathcal{B}_{p,p}^{s}(\Omega)}$. Here, the implicit constant depends only on $\Gamma$, $s$ and $p$.
@@ 5.1 The conjugate operator
Assume that $\Omega$ is a bounded chord-arc domain and $\mathbb{D}$ is the unit disk. Let $z_{0}$ be a point in $\Omega$. If $u$ is harmonic in $\Omega$ it is well known that there exists a unique harmonic function $\tilde{u}$ in $\Omega$ such that $\tilde{u}(z_{0})=0$ and $u+i\tilde{u}$ is holomorphic in $\Omega$. In the case of $\Omega=\mathbb{D}$, $\tilde{u}$ is the harmonic extension (Poisson) of $H(f)$ where $H$ is the Hilbert transform:
@@
By the Cauchy-Riemann equations, $|\nabla\tilde{u}|=|\nabla u|$. Thus, we see that $\mathcal{B}_{p,p}^{s}(\Omega)$ is conjugate-invariant. We denote the subspace of $\mathcal{B}_{p,p}^{s}(\Omega)$ consisting of holomorphic functions by $\mathcal{HB}_{p,p}^{s}(\Omega)$, i.e., the set of $u+i\tilde{u}$ with $u,\tilde{u}\in\mathcal{B}_{p,p}^{s}(\Omega)$.
A necessary condition for $B_{p,p}^{s}(\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega)$ to hold is thus that $B_{p,p}^{s}(\Gamma)$ is stable by conjugation. Let $\varphi:\,\mathbb{D}\to\Omega$ be a Riemann map with $\varphi(0)=z_{0}$. Recall that $t\mapsto z(t)$ denotes the arc-length parametrization of $\Gamma$ and $\lambda(e^{it})=z(t)$ so that $\varphi=\lambda\circ h$ where $h$ is an absolutely continuous homeomorphism of $\mathbb{T}$ with $|h^{\prime}|=|\varphi^{\prime}|$. Recall that $\lambda$ is a bi-Lipschitz homeomorphism from $\mathbb{T}$ onto $\Gamma$. A simple computation shows that if $f\in B_{p,p}^{s}(\Gamma)$; that is, $g=f\circ\lambda\in B_{p,p}^{s}(\mathbb{T})$ and $u$ is its harmonic extension to $\Omega$, then $\tilde{u}$ is the harmonic extension of $\tilde{f}=\tilde{g}\circ\lambda^{-1}$. Here,
@@
It is known (see [9], p.247 in [16]) that $V_{h}^{-1}HV_{h}$ is bounded on $L^{p}(\mathbb{T})$, $1<p<\infty$, if and only if $|h^{\prime}|=|\varphi^{\prime}|$ belongs to $A_{p}$ on $\mathbb{T}$. Based on this comment on the limiting case $s=0$, we now state a theorem about the general case $0\leq s\leq 1$.
@@
For $p>1,\,0\leq s\leq 1$, suppose that $\Gamma$ is such that
@@
Then $B_{p,p}^{s}(\Gamma)$ is stable by conjugation, and, moreover, the conjugate operator $f\mapsto\tilde{f}$ on $B_{p,p}^{s}(\Gamma)$ is bounded.
By this theorem, we define $\text{HB}_{p,p}^{s}(\Gamma)$ as the subspace of $B_{p,p}^{s}(\Gamma)$ consisting of functions $f+i\tilde{f}$ with $f,\tilde{f}\in B_{p,p}^{s}(\Gamma)$, and $\text{HB}_{p,p}^{s}(\Omega)$ as the analytic subspace of $B_{p,p}^{s}(\Omega)$.
Before going to the proof, let us comment on the hypotheses of Theorem 5.1. If $\omega$ is an $A_{\infty}$-weight, let us define $\delta(\omega)=\inf\{q>1:\,\omega\in A_{q}\}$. It is known [9] that if $\omega\in A_{q}$ for some $q>1$ then $\delta(\omega)<q$ (see p.254 in [16]). The above theorem means that given $p>1$ and the class of curves $\Gamma$ for which the claim is true depends on $p$. If we reverse the point of view and start with a curve $\Gamma$ satisfying $|\varphi^{\prime}|\in A_{2}$, then by defining $\delta=\delta(|\varphi^{\prime}|)$, the claim is true for
$\delta<p<\frac{\delta}{\delta-1}.$
On the other hand, it is known (see [21]) that $\Gamma$ being chord-arc implies $|\varphi^{\prime}|$ having $A_{\infty}$ but there are examples of chord-arc curves such that $|\varphi^{\prime}|\notin A_{2}$ (see [23]), we thus know that the chord-arc condition is not sufficient for Theorem 5.1 to hold.
@@
Specially, a natural condition on the curve $\Gamma$ implying that $|\varphi^{\prime}|\in A_{2}$ is the radial-Lipschitz condition. Indeed, if $\Gamma$ is a radial-Lipschitz curve, then it can be shown that $|\varphi^{\prime}|$ satisfies the Helson-Szegö condition (see subsection 5.3 for a specific proof): $\log|\varphi^{\prime}|=u+Hv$ with $u\in L^{\infty},\;v\in L^{\infty}$ and $\|v\|_{\infty}<\pi/2$, which is equivalent to $|\varphi^{\prime}|\in A_{2}$ as follows from the Helson-Szegö theorem ([16]). We may now state:
@@
For $0\leq s\leq 1$, if $\Gamma$ is such that $|\varphi^{\prime}|\in A_{2}$ then $B_{2,2}^{s}(\Gamma)$ is stable by conjugation, and moreover, the conjugate operator $f\mapsto\tilde{f}$ in $B_{2,2}^{s}(\Gamma)$ is bounded. The conclusion holds in particular for radial-Lipschitz curves.
@@ 5.2 The operator $V_{s,p}$ on $\mathcal{B}^{s}_{p,p}(\Omega)$
Let $\Omega,\,\Omega^{\prime}$ be two Jordan domains containing 0 and $\varphi$ a holomorphic diffeomorphism from $\Omega$ onto $\Omega^{\prime}$ fixing 0. Let $f\in\mathcal{HB}_{p,p}^{s}(\Omega^{\prime})$. Using the change of variables $\zeta=\varphi(z)$ we have
@@
By the Koebe distortion theorem (see (2.1)) there exists a universal constant $C>1$ such that
@@
Notice that for s=11/ps=1-1/p, this operator is nothing more than the composition by $\varphi$ and that, in this case, $V_{1-1/p,p}$ is an isometry.
For convenience of later use, when $p=2$, the operator $V_{s,p}$ will be simplified to $V_{s}$.
In order to better understand the operator $V_{s}$ let us rewrite it by using an integration by parts:
@@
where the supremum is taken over all sub-arcs $I$ of $\mathbb{T}$ and $g_{I}$ denotes the integral mean of $g$ over $I$. It is known that $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$ if the domain $\varphi(\mathbb{D})$ is chord-arc, but not vise versa. By Fefferman-Stein, $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$ if and only if
@@
is a Carleson measure in $\mathbb{D}$ (see, e.g., Ch.VI in [16]).
@@
Suppose $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$; that is, $\mu$ is a Carleson measure in $\mathbb{D}$. Let $f\in E^{2}(\Omega)$ be such that $T_{0}f\in E^{2}(\mathbb{D})$. Then by Carleson (see, e.g., Theorem 3.9 in [16])
@@
Combined with (5.2), this completes the proof of the statement $(1)$.
@@
or, in other words, $\mathsf{g}(f\circ\varphi)\in L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$. Since $\Gamma$ is assumed to be chord-arc, we have that $|\varphi^{\prime}|$ has the weight $A_{\infty}$ on $\mathbb{T}$. We can then apply a theorem of Gundy-Wheeden [19] (see also [21]) which implies that $\mathsf{g}(f\circ\varphi)\in L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$ if and only if the non-tangential maximal function $\mathsf{n}(f\circ\varphi)$ of $f\circ\varphi$ is in $L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$. On the other hand, for any "circular curves" $\Gamma_{r}$ we have that (see p.233 in [21])
@@
Notice that for this theorem, we do not need $|\varphi^{\prime}|$ having $A_{2}$ on $\mathbb{T}$. This condition is nevertheless necessary for the following corollary to hold by Corollary 5.2 since if it is not attached then the (real) space of real parts of $E^{2}(\Omega\!\to\!\Gamma)$-functions is a proper subspace of the real space $L_{\mathbb{R}}^{2}(\Gamma,d\sigma)$
@@
Let $\Omega$ be a chord-arc domain bounded by $\Gamma$. If $|\varphi^{\prime}|\in A_{2}$ then $B_{2,2}^{0}(\Gamma)=\mathcal{B}_{2,2}^{0}(\Omega)$.
5.3 The equality $\mathcal{B}_{2,2}^{s}(\Omega)=B_{2,2}^{s}(\Gamma)$
In this subsection, our goal is to prove the following theorem, which is one of the main theorems in this paper.